# Mass-imbalance effect on the cluster formation in a one-dimensional Fermi gas with coexistent $s$- and $p$-wave interactions

### v1

@@ III.2 In-medium case
Figure 4: Contour plot of in-medium three-body energy $E_{3}^{aab}/E_{0}$ in a plane of mass $m_{a}/m_{0}$ and $m_{b}/m_{0}$. The mass $m_{0}$ and energy $E_{0}=k_{\rm F}^{2}/(2m_{0})$ are taken as the reference scales. The interaction strengths are fixed at $1/(k_{\rm F}a_{s})=1/(k_{\rm F}a_{p})=1.5$. The momentum cutoff is taken as $10k_{\rm F}$.
@@
To see effects from the mass imbalance to the in-medium three-body ground-state energy, in Fig. 4, we plot the contour plot of in-medium three-body energy $E_{3}^{aab}/E_{0}$ in a plane of mass $m_{a}/m_{0}$ and $m_{b}/m_{0}$. For comparison, the mass $m_{0}$ and energy $E_{0}=k_{\rm F}^{2}/(2m_{0})$ are taken as the reference scales. The interaction strengths are fixed at $1/(k_{\rm F}a_{s})=1/(k_{\rm F}a_{p})=1.5$. The momentum cutoff $\Lambda$ is taken as $10k_{\rm F}$. In the present two-component Fermi gas, as for one species of atom with a fixed mass, if we gradually enlarge the mass of another one, the absolute value of $E_{3}^{aab}$ monotonically decreases. Such a qualitative behavior is a result of fixed interaction strengths. Moreover, it is also seen that a smaller reduced mass $m_{r}=m_{a}m_{b}/(m_{a}+m_{b})$ gives an negatively larger $E_{3}^{aab}$ in Fig. 4.
@@
Figure 5: Lower-lying excited state of Cooper triple phases with different configurations in a plane of two-body $p$-wave interaction strengths. The mass ratio is taken as $m_{a}=2m_{b}$, and the momentum cutoffs are taken as $\Lambda/k_{\rm F}=10$.
@@
In order to see the competition between the in-medium three-body correlations with different configurations in the present system, by fixing the mass ratio as $m_{a}=2m_{b}$, we plot the lower-lying excited state of Cooper triple phases with different configurations in a plane of two-body $p$-wave interaction strengths in Fig. 5. Here the dimensionless $p$-wave coupling strengths are defined as $\tilde{U}_{i}=2m_{r}k_{\rm F}U_{i}$. In this paper, we consider the attractive two-body $p$-wave interactions, which indicates that a larger absolute value of $\tilde{U}_{i}$ represents a stronger attraction. At a fixed $p$-wave interaction strength $\tilde{U}_{a}$ among atoms $a$, when the $p$-wave interaction strength $\tilde{U}_{b}$ among atoms $b$ increases to a certain value, the dominant (lower-lying) Cooper triple state changes from the aabaab configuration to the abbabb one. The similar behavior holds for a fixed $p$-wave interaction strength $\tilde{U}_{b}$ among atoms $b$. The competition between two different configurations of Cooper triples is similar to the in-vacuum case shown in Fig. 3, which results from the mass difference.
Figure 6: Phase diagram of $s$-wave pairing ($|E_{2,s}|>|E_{2,p}^{ii}|,\ |E_{3}^{iij}|$), $p$-wave pairings ($|E_{2,p}^{ii}|>|E_{2,s}|,\ |E_{3}^{iij}|$), and Cooper tripling ($|E_{3}^{iij}|>|E_{2,s}|,\ |E_{2,p}^{ii}|$) in the plane of $1/(k_{\rm F}a_{s})$ and $1/(k_{\rm F}a_{p})$ ($i,j=a,b$). The mass ratio is taken as $m_{a}=2m_{b}$, and the momentum cutoffs are taken as $\Lambda/k_{\rm F}=10$.
@@
At last, by taking the mass ratio as $m_{a}=2m_{b}$, we summarize the ground-state phase diagram of $s$- and $p$-wave Cooper pairing states, and the Cooper triple states with different configurations in Fig. 6. The boundary lines between different kinds of Cooper clustering phases are determined in such a way that the one between tripling (iijiij configuration) and $s$-wave pairing, that between tripling (iijiij configuration) and $p$-wave pairing ($ii$ configuration), and that between $s$- and $p$-wave pairing ($ii$ configuration) are given by $E_{3}^{iij}=E_{2,s}$, $E_{3}^{iij}=E_{2,p}^{ii}$, and $E_{2,s}=E_{2,p}^{ii}$, respectively. The range of scattering parameters in Fig. 6 $0.5\leq 1/(k_{\rm F}a_{s})\leq 3$ and $1.5\leq 1/(k_{\rm F}a_{p})\leq 1.5$] are adopted as same one in Ref. [5], which are also similar values were realized in a recent experimental work [47]. Such a phase diagram captures interesting features associated with competing $s$- and $p$-wave pairings and moreover tripling. In order to investigate the physics behind more, we also plot in-medium clustering energies as functions of scattering lengths in Fig. 2. From Fig. 2, it is shown that the presence of the Fermi sea reshapes the formation of pairing and three-body states. Moreover, by comparing Fig. 2(a) [or (b)] and (c), the presence of Fermi surface affects three-body energies much more than pairing ones. Consequently, it is seen that when $s$- or $p$-wave interaction strength is comparably weak, the pairing phases survive, which is different to the in-vacuum case. As for the pairing states, the $p$-wave pair ($aa$ configuration) phase is always suppressed due to the mismatch of Fermi energies between atoms $a$ and $b$, where the Cooper instability among heavier atoms becomes weaker. As for the tripling states, the Cooper triple (aabaab configuration) phase survives due to the coexistence of both $s$- and $p$-wave interactions, where the effect from the mismatch of Fermi energies is partially offset by the inclusion of $s$-wave attraction. In the meanwhile, similar to the in-vacuum case shown in Fig. 3, there is a competition between Cooper triple (aabaab configuration) phase and Cooper triple (abbabb configuration) phase. On the one hand, we have numerically checked that the qualitative behaviors of the boundary line between $s$-wave pair phase and Cooper triple (aabaab configuration) phase, and that between Cooper triple (aabaab configuration) phase and Cooper triple (abbabb configuration) phase are unchanged even up to a extremely large $s$-wave interaction strength [such as $1/(k_{\rm F}a_{s})=10.0$]. It indicates that the competition between different configurations of Cooper triple phases may also result from the difference of scatterings between $s$-wave dimer ($ab$ configuration) and each component of atoms. In the meanwhile, at a fixed $s$-wave interaction strength, $bb$ configuration with a smaller mass also becomes more tightly bound than the $aa$ configuration as the $p$-wave scattering length increases due to the mismatch of Fermi energy between atoms $a$ and $b$. In other words, the Cooper triple (abbabb configuration) is expected to be stabler than the Cooper triple (aabaab configuration) in the stronger $p$-wave interaction regime. As for the cutoff dependence, we have numerically checked that while the regime of Cooper triple phases gets enlarged with larger momentum cutoff $\Lambda$, the boundaries between different kinds of pairings remain robust. Such a cutoff dependenc is also similar to that in the previous work [5]. Even though the specific value of momentum cutoff in the practical calculations is needed to compare our results with the experiments, our phase diagram would be useful to understand the qualitative features of hybridized $s$- and $p$-wave interacting systems with mass imbalance and different configurations of clustering.
@@ IV Summary and perspectives
As a step further of previous work [5], in this paper we have investigated the mass imbalance and different kinds of clustering configurations in the system with the coexistence $s$- and $p$-wave interactions. We have investigated both the in-vacuum and in-medium two- and three-body correlations in the mass-imbalanced two-component Fermi gas. The solutions of in total five kinds of stable in-medium two- and three-body cluster states have been found in this system by solving the variational equations. By taking a certain mass ratio between two components of fermions, we have featured out a ground-state phase diagram consisting of $s$- and $p$-wave pairing states, and the Cooper triple states with different configurations in a plane of $s$- and $p$-wave scattering lengths.

### v2

@@ III.2 In-medium case
Figure 4: Contour plot of the in-medium three-body energy $E^{aab}_{3}/E_{0}$ as a function of the masses $m_{a}/m_{0}$ and $m_{b}/m_{0}$. Here $E_{0}=k_{F}^{2}/(2m_{0})$ is the reference energy scale. The energy is measured relative to the filled Fermi sea of noninteracting fermions, and a more negative value indicates a more strongly bound in-medium trimer. The interaction strengths are fixed to $1/(k_{F}a_{s})=1/(k_{F}a_{p})=1.5$, and the momentum cutoff is $\Lambda/k_{F}=10$.
@@
To see effects from the mass imbalance to the in-medium three-body energy, in Fig. 4, we plot the contour plot of in-medium three-body energy $E_{3}^{aab}/E_{0}$ in a plane of mass $m_{a}/m_{0}$ and $m_{b}/m_{0}$. For comparison, the mass $m_{0}$ and energy $E_{0}=k_{\rm F}^{2}/(2m_{0})$ are taken as the reference scales. The interaction strengths are fixed at $1/(k_{\rm F}a_{s})=1/(k_{\rm F}a_{p})=1.5$. The momentum cutoff $\Lambda$ is taken as $10k_{\rm F}$. In the present two-component Fermi gas, as for one species of atom with a fixed mass, if we gradually enlarge the mass of another one, the absolute value of $E_{3}^{aab}$ monotonically decreases. Such a qualitative behavior is a result of fixed interaction strengths. Moreover, it is also seen that a smaller reduced mass $m_{r}=m_{a}m_{b}/(m_{a}+m_{b})$ gives a more negative $E_{3}^{aab}$ in Fig. 4.
@@
Figure 5: Lower-lying excited state of Cooper trimer phases with different configurations in a plane of two-body $p$-wave interaction strengths. The mass ratio is taken as $m_{a}=2m_{b}$, and the momentum cutoffs are taken as $\Lambda/k_{\rm F}=10$.
@@
In order to see the competition between the in-medium three-body correlations with different configurations in the present system, by fixing the mass ratio as $m_{a}=2m_{b}$, we plot the lower-lying excited state of Cooper trimer phases with different configurations in a plane of two-body $p$-wave interaction strengths in Fig. 5. Here the dimensionless $p$-wave coupling strengths are defined as $\tilde{U}_{i}=2m_{r}k_{\rm F}U_{i}$. In this paper, we consider the attractive two-body $p$-wave interactions, which indicates that a larger absolute value of $\tilde{U}_{i}$ represents a stronger attraction. At a fixed $p$-wave interaction strength $\tilde{U}_{a}$ among atoms $a$, when the $p$-wave interaction strength $\tilde{U}_{b}$ among atoms $b$ increases to a certain value, the dominant (lower-lying) Cooper trimer state changes from the aabaab configuration to the abbabb one. The similar behavior holds for a fixed $p$-wave interaction strength $\tilde{U}_{b}$ among atoms $b$. The competition between two different configurations of Cooper trimers is similar to the in-vacuum case shown in Fig. 3, which results from the mass difference.
Figure 6: Phase diagram of $s$-wave pairing, $p$-wave pairing, and Cooper trimer states in the plane of $1/(k_{F}a_{s})$ and $1/(k_{F}a_{p})$. The phase boundaries are determined by comparing the in-medium cluster energies, with the dominant instability corresponding to the most negative energy. All energies are measured relative to the energy of the filled Fermi sea of noninteracting fermions. The mass ratio is fixed to $m_{a}/m_{b}=2$, and the momentum cutoff is $\Lambda/k_{F}=10$.
@@
At last, by taking the mass ratio as $m_{a}=2m_{b}$, we summarize the resulting hierarchy of dominant instabilities of $s$- and $p$-wave Cooper pairing states, and the Cooper trimer states with different configurations in Fig. 6. The boundary lines between different kinds of Cooper clustering phases are determined in such a way that the one between Cooper trimer state (iijiij configuration) and $s$-wave pairing, that between Cooper trimer state (iijiij configuration) and $p$-wave pairing ($ii$ configuration), and that between $s$- and $p$-wave pairing ($ii$ configuration) are given by $E_{3}^{iij}=E_{2,s}$, $E_{3}^{iij}=E_{2,p}^{ii}$, and $E_{2,s}=E_{2,p}^{ii}$, respectively. The range of scattering parameters in Fig. 6 $0.5\leq 1/(k_{\rm F}a_{s})\leq 3$ and $1.5\leq 1/(k_{\rm F}a_{p})\leq 1.5$] are adopted as same one in Ref. [5], which are also similar values were realized in a recent experimental work [47]. Such a phase diagram captures interesting features associated with competing $s$- and $p$-wave pairings and moreover trimer formation. In order to further investigate the underlying physics, we also plot in-medium clustering energies as functions of scattering lengths in Fig. 2. From Fig. 2, it is shown that the presence of the Fermi sea reshapes the formation of pairing and three-body states. Moreover, by comparing Fig. 2(a) [or (b)] and (c), the presence of Fermi surface affects three-body energies much more than pairing ones. Consequently, it is seen that when $s$- or $p$-wave interaction strength is comparably weak, the pairing phases survive, which is different to the in-vacuum case. As for the pairing states, the $p$-wave pair ($aa$ configuration) phase is always suppressed due to the mismatch of Fermi energies between atoms $a$ and $b$, where the Cooper instability among heavier atoms becomes weaker. As for the trimer states, the Cooper trimer (aabaab configuration) phase survives due to the coexistence of both $s$- and $p$-wave interactions, where the effect from the mismatch of Fermi energies is partially offset by the inclusion of $s$-wave attraction. In the meanwhile, similar to the in-vacuum case shown in Fig. 3, there is a competition between Cooper trimer (aabaab configuration) phase and Cooper trimer (abbabb configuration) phase. On the one hand, we have numerically checked that the qualitative behaviors of the boundary line between $s$-wave pair phase and Cooper trimer (aabaab configuration) phase, and that between Cooper trimer (aabaab configuration) phase and Cooper trimer (abbabb configuration) phase are unchanged even up to a extremely large $s$-wave interaction strength [such as $1/(k_{\rm F}a_{s})=10.0$]. It indicates that the competition between different configurations of Cooper trimer phases may also result from the difference of scatterings between $s$-wave dimer ($ab$ configuration) and each component of atoms. Meanwhile, at a fixed $s$-wave interaction strength, $bb$ configuration with a smaller mass also becomes more tightly bound than the $aa$ configuration as the $p$-wave scattering length increases due to the mismatch of Fermi energy between atoms $a$ and $b$. In other words, the Cooper trimer (abbabb configuration) is expected to be more stable than the Cooper trimer (aabaab configuration) in the stronger $p$-wave interaction regime.
Table 1: Cutoff dependence of boundary point $1/(k_{\rm F}a_{p})$ of $s$-wave pairing and Cooper trimer state (aabaab configuration).
$1/(k_{\rm F}a_{s})$ $\Lambda/k_{\rm F}=8$ $\Lambda/k_{\rm F}=10$ $\Lambda/k_{\rm F}=12$
$2.8$ $0.021$ 0.386-0.386 0.522-0.522
$2.9$ 0.001-0.001 0.450-0.450 0.666-0.666
$3.0$ 0.175-0.175 0.500-0.500 0.850-0.850
We have numerically checked the cutoff dependence, and the physical two-body scattering parameters $a_{s}$ and $a_{p}$ are kept fixed when varying the momentum cutoff $\Lambda$. As shown in Table 1, by scanning several boundary points of $s$-wave pairing and Cooper trimer state (aabaab configuration), we find that increasing $\Lambda$ enlarges the parameter region where the Cooper trimer instability dominates. While the quantitative locations of the phase boundaries exhibit moderate cutoff-dependent shifts, the qualitative structure of the phase diagram remains robust, including the hierarchy between pairing and Cooper trimer instabilities and the competition between different Cooper trimer configurations. The momentum cutoff is introduced as an ultraviolet regularization for the zero-range interactions. Even though the same cutoff corresponds to different single-particle energy scales for the two species in the mass-imbalanced case, the low-energy physics discussed here is governed by the relative momentum and remains unaffected. Such a cutoff dependence is also similar to that in the previous work [5]. While the specific value of momentum cutoff in the practical calculations is needed to compare our results with the experiments, our phase diagram would be useful to understand the qualitative features of hybridized $s$- and $p$-wave interacting systems with mass imbalance and different configurations of clustering.
We note that effects beyond the present variational treatment, such as medium polarization, self-energy dressing, and finite lifetime of Cooper clusters, are not included. These effects may quantitatively renormalize the cluster energies and phase boundaries. Nevertheless, as long as the system is in the regime where pairing or clustering instabilities emerge continuously from the Fermi surface, the relative ordering of instabilities is expected to remain qualitatively reliable. Furthermore, the possible emergence of four-body bound states, such as a tetramer in aabbaabb configuration, cannot be excluded in principle, especially in the presence of both $a$-$a$ and $b$-$b$ odd-parity interactions. The explicit treatment of such four-body correlations is beyond the scope of the present work and is left for future studies.
Before drawing the conclusion, it is noted that although most results are presented for a representative mass ratio $m_{a}/m_{b}=2$, we have verified that the qualitative features of the pairing–trimer competition and the structure of the phase diagram persist for other mass ratios within a reasonable range. In particular, varying the mass ratio mainly leads to quantitative shifts of phase boundaries, while the hierarchy between pairing and Cooper trimer formation remains unchanged.
@@ IV Summary and perspectives
As a step further of previous work [5], in this paper we have investigated the mass imbalance and different kinds of clustering configurations in the system with the coexistence $s$- and $p$-wave interactions. We have investigated both the in-vacuum and in-medium two- and three-body correlations in the mass-imbalanced two-component Fermi gas. The solutions of in total five kinds of stable in-medium two- and three-body cluster states have been found in this system by solving the variational equations. By taking a certain mass ratio between two components of fermions, we have mapped out a lowest-energy cluster-instability diagram consisting of $s$- and $p$-wave pairing states, and the Cooper trimer states with different configurations in a plane of $s$- and $p$-wave scattering lengths.